OpenStudy (anonymous):

I'm confused again... Instead of the y = mx + b I'm given no formula just told to write the equation for the following relation. which is: B = {(x,y): (2,3), (4,4), (6,5)} The answer is 2y = x + 4 I tried the y = mx + b formula, which turn to be - (4-3)/(4-2)= 1÷2=2 (5-4)/(6-4)= 1÷2=2 (5-3)/(6-2)= 2÷4=2 now its y = 2x + b 3=2(2) + b -4 + 3 = b -1 = b It would all be y = 2x -1 but since i know the answer its not that, is it a different formula and if so how do you know when to use the right formula when they don't tell you which one?

OpenStudy (mathstudent55):

I think you mean "write the formula for the following _relation_."

OpenStudy (mathstudent55):

We can still use y = mx + b First, we use any pair of 2 of the given points to find the slope. Do you know how to find the slope of a line given two points on the line?

OpenStudy (mathstudent55):

The slope, \(m\), of the line that passes through points \((x_1, y_1)\) and \((x_2, y_2)\) is \(m = \dfrac{y_2 - y_1}{x _2 - x_1} \)

OpenStudy (anonymous):

first find m= (y-y0)/(x-x0) choice any 2 points

OpenStudy (anonymous):

then use y-yo=m(x-x0) use any points

OpenStudy (anonymous):

i used 1 and 2 points y=1/2x+2

OpenStudy (anonymous):

multply this by 2 2y=x+4

OpenStudy (mathstudent55):

Once you find the slope, plug it in the equation y = mx + b Then use any point and enter its coordinates for x and y in the equation above (also using the value of m you obtained above), and solve for b. Then plug in b into the equation.

OpenStudy (anonymous):

so its \[M = \frac{ 3-4 }{ 2-4 }= \frac{ 1 }{ 2 }\] I'm stuck on y-yo=m(x-x0) would it be written 3-4 = m(2-4) or do I add that 0 by x0 to my 4 making it 40? but isn't the slope 1/2 so I'd go to that y = mx+b I end up with 3 = 2(2) + b -4+3=b 1/2-1=b -0.5=b which isn't right...

OpenStudy (mathstudent55):

No. The idea of y - y0 = m(x - x0) is to leave x and y as they are, and use the point's coordinates in x0 and y0. Like this: y - 3 = (1/2)(x - 2)

OpenStudy (mathstudent55):

The above was using the correct slope you found of 1/2 and the point (2, 3)

OpenStudy (anonymous):

So now I'd bring in the y = mx + b? Making it... y - 3 = (1/2)(2) + b y -3=1+b -1-3=b 2=b then it'd be 2y = x... and where's the 4... or is it y = 1/2x + 2(2) y=1/2x + 4 then we do y-3=(1/2)(2) + 4 -1-3=4 2=4 then write it out like 2y = x + 4... and if so, do I leave x as it is, from the y - y0 = m(x - x0) formula since one x was left alone and the same with 2y since one y was left alone?

OpenStudy (mathstudent55):

There are several ways of doing this problem. I'll show you two ways. 1) Using \(y = mx + b\) 2) Using \(y - y_0 = m(x - x_0) \) For both ways you first need to find m, the slope, so let's start with that. I am using the first two given points, (2, 3) and (4, 4).: \(m = \dfrac{y_2 - y_1}{x_2 - x_1} = \dfrac{4 - 3}{4 - 2} = \dfrac{1}{2}\) You already had the slope above that you calculated correctly. Now let's follow method 1) y = mx + b First, we plug in the slope into the equation y = mx + b to get \(y = \dfrac{1}{2}x + b\) We need to find b. We use any point of the three given points in y = mx + b and solve for b. Let's use point (2, 3) \(3 = \dfrac{1}{2}\times 2 + b\) \(3 = 1 + b\) \(2 = b\) Now that we have b = 2, we substitute this value into \(y = \dfrac{1}{2}x + b\) to get: \(y = \dfrac{1}{2}x + 2\) This is the correct equation in slope-intercept form. Now we multiply both sides by 2 to get: \(2y = x + 4\) which is the answer you know is correct.

OpenStudy (mathstudent55):

Now let's use method 2) \(y - y_0 = m(x - x_0)\) Here we need the slope, m, which we already have, and we need a point. The coordinates of the point we choose to use are substituted into \(x_0\) and \(y_0\). \(y - y_0 = m(x - x_0)\) Let's use the last given point, (6, 5) \(y - 5 = \dfrac{1}{2}(x - 6) \) Multiply both sides by 2: \(2y - 10 = x - 6\) Add 10 to both sides: \(2y = x + 4\) Once again, as expected, we find the same answer.

OpenStudy (mathstudent55):

This is how you solve this problem using these two methods.

OpenStudy (mathstudent55):

In your last response, you were using y = mx + b. You did calculate b correctly as b = 2. Then you did not know how to apply the b = 2. I hope the explanation above clears this up. If you still have any questions, feel free to ask.

OpenStudy (anonymous):

I think I have it, I'll test it out on a few other problems to make sure I got it, and thank you for the help.

OpenStudy (mathstudent55):

You're welcome.